# The Invisible Hand of AI Libraries Shaping Open Source Projects and Communities

### v1

@@ IV-A5 Complexity and Maintainability ($RQ_{3}$)
To investigate how adopting AI libraries influences the internal structure and maintainability of open-source projects at the code level ($RQ_{3}$), we will collect SMs using SciTools Understand, a widely-used static code analysis tool [scitools_understand_2025].
@@ IV-B2 Community Engagement & Development Process ($RQ_{2}$)
We tested $H_{\mathcal{NRM}}$ with the Anderson-Darling (AD) test [andersondarling1952]. The AD test assesses whether data samples derive from a specific probability distribution, such as the normal distribution. AD measures the difference between the sample data and the expected values from the tested distribution. More specifically, it evaluates differences in the cumulative distribution function (CDF) between the observed data and the hypothesized distribution [andersondarling1952].
@@
The AD and the Shapiro-Wilk (SW) [shaphiro1965analysis] are both statistical tests used to assess the normality of data. The SW test focuses on the correlation between the observed data and the expected values under a normal distribution, emphasizing the smallest and largest values in the dataset. Therefore, according to [mishra2019descriptive], SW is more appropriate when targeting small datasets ($\leq 50$ samples). On the other hand, the AD test considers a broader range of values, including those in the middle of the distribution, providing a more sensitive evaluation of normality, especially for larger sample sizes than our own. Therefore, we preferred AD over SW [Stephens1974EDF]. AD is also considered one of the most powerful statistical tools for detecting most departures from normality [Stephens1974EDF, robredo_manero_analyzing_2024, esposito_call_2025].
@@
In case of acceptance of the null hypotheses ($H_{\mathcal{NRM}}$), similarly to the normality testing sample concerns, we will test $H_{1(3\rightarrow 5)}$ with the Z-statistic [sprinthall2011basic]. We can use the Z-statistic to assess the significance of observed differences between sample data and population parameters. Calculated by standardizing sample statistics based on known population parameters, the Z-statistic quantifies the extent to which observed results deviate from what would be expected under the null hypothesis [sprinthall2011basic, casella2002statistical, montgomery2020applied].
@@
Interpreting the Z-statistic involves comparing its value with critical values of the standard normal distribution, typically determined by the chosen significance level [casella2002statistical]. If the calculated Z statistic falls within the critical region, it suggests that the observed results are extreme enough to reject the null hypothesis in favor of the alternative [montgomery2020applied].
@@
In case of rejection of the null hypothesis ($H_{\mathcal{N}}$), we will test $H_{1(3\rightarrow 5)}$ using the Wilcoxon signed rank test (WT) [pub.1102728208], which is a nonparametric statistical test that compares two related samples or paired data. WT uses the absolute difference between the two observations to classify and then compare the sum of the positive and negative differences. The test statistic is the lowest of both. We selected WT to test $H_{1(3\rightarrow 8)}$ because RMs’ observations would not be normally distributed; hence, we used it instead of the paired t-test, which assumes a normal data distribution.
@@
We set our alpha to 0.01. We decided to reduce the alpha value from the standard value of 0.05 due to the many statistical tests we will perform and therefore to reach a good balance between Type I and II error [abac.12172].
@@ IV-B3 Complexity and Maintainability ($RQ_{3}$)
Similar to $RQ_{2}$, in the case of acceptance of the null hypothesis ($H_{0\mathcal{N}SM}$), we will test the hypothesis $H_{1(6\rightarrow 8)}$ using the Z-statistic [sprinthall2011basic].
However, in the case of rejection of the null hypothesis ($H_{0\mathcal{N}SM}$), we will instead adopt the Wilcoxon signed-rank test (WT) [pub.1102728208] to evaluate $H_{1(6\rightarrow 8)}$.
@@ IV-C Replicability and Verifiability
Our work will adhere to the Open Science principles[open_science_principles]. We will prepare the replication package, containing raw data and scripts, and share it on Zenodo, according to FAIR principles [force11_fair], i.e., Findable, Accessible, Interoperable, and Reusable to foster applicability and verifiability and bolster the feasibility of future works.
@@ V Threats to Validity
In this section, we discuss the threats to the validity of our study. We categorized the threats into Construct, Internal, External, and Conclusion, following the guidelines defined by Wohlin et al. [DBLP:books/daglib/0029933].
Construct Validity. Our specific design choices, including our measurement process and data filtering, may impact our results. To address this threat, we based our choice on past studies [dabic2021sampling, tufano2024unveiling, 10.1145/3453478, DBLP:journals/tse/YuBSM21] and used well-established guidelines in designing our methodology [DBLP:journals/ese/RunesonH09, Basili1994].
Internal Validity. Our study relies on a large-scale analysis of 6,323 repositories, which can potentially be biased from the project selection. We address this threat by designing our inclusion criteria based on past studies [10.1145/3453478, DBLP:journals/tse/YuBSM21, dabic2021sampling].
@@ Acknowledgments
This work has been funded by the Research Council of Finland (grants n. 359861 and 349488 - MuFAno) and Business Finland (grant 6GSoft[akbar_6gsoft_2024]).

### v2

@@ IV-A5 Complexity and Maintainability ($RQ_{3}$)
To investigate how adopting AI libraries influences the internal structure and maintainability of open-source projects at the code level ($RQ_{3}$), we will collect SMs using SciTools Understand, a widely-used static code analysis tool [58].
@@ IV-B2 Community Engagement & Development Process ($RQ_{2}$)
We tested $H_{\mathcal{NRM}}$ with the Anderson-Darling (AD) test [5]. The AD test assesses whether data samples derive from a specific probability distribution, such as the normal distribution. AD measures the difference between the sample data and the expected values from the tested distribution. More specifically, it evaluates differences in the cumulative distribution function (CDF) between the observed data and the hypothesized distribution [5].
@@
The AD and the Shapiro-Wilk (SW) [59] are both statistical tests used to assess the normality of data. The SW test focuses on the correlation between the observed data and the expected values under a normal distribution, emphasizing the smallest and largest values in the dataset. Therefore, according to [49], SW is more appropriate when targeting small datasets ($\leq 50$ samples). On the other hand, the AD test considers a broader range of values, including those in the middle of the distribution, providing a more sensitive evaluation of normality, especially for larger sample sizes than our own. Therefore, we preferred AD over SW [62]. AD is also considered one of the most powerful statistical tools for detecting most departures from normality [62, 56, 32].
@@
In case of acceptance of the null hypotheses ($H_{\mathcal{NRM}}$), similarly to the normality testing sample concerns, we will test $H_{1(3\rightarrow 5)}$ with the Z-statistic [60]. We can use the Z-statistic to assess the significance of observed differences between sample data and population parameters. Calculated by standardizing sample statistics based on known population parameters, the Z-statistic quantifies the extent to which observed results deviate from what would be expected under the null hypothesis [60, 13, 51].
@@
Interpreting the Z-statistic involves comparing its value with critical values of the standard normal distribution, typically determined by the chosen significance level [13]. If the calculated Z statistic falls within the critical region, it suggests that the observed results are extreme enough to reject the null hypothesis in favor of the alternative [51].
@@
In case of rejection of the null hypothesis ($H_{\mathcal{N}}$), we will test $H_{1(3\rightarrow 5)}$ using the Wilcoxon signed rank test (WT) [68], which is a nonparametric statistical test that compares two related samples or paired data. WT uses the absolute difference between the two observations to classify and then compare the sum of the positive and negative differences. The test statistic is the lowest of both. We selected WT to test $H_{1(3\rightarrow 8)}$ because RMs’ observations would not be normally distributed; hence, we used it instead of the paired t-test, which assumes a normal data distribution.
@@
We set our alpha to 0.01. We decided to reduce the alpha value from the standard value of 0.05 due to the many statistical tests we will perform and therefore to reach a good balance between Type I and II error [43].
@@ IV-B3 Complexity and Maintainability ($RQ_{3}$)
Similar to $RQ_{2}$, in the case of acceptance of the null hypothesis ($H_{0\mathcal{N}SM}$), we will test the hypothesis $H_{1(6\rightarrow 8)}$ using the Z-statistic [60].
However, in the case of rejection of the null hypothesis ($H_{0\mathcal{N}SM}$), we will instead adopt the Wilcoxon signed-rank test (WT) [68] to evaluate $H_{1(6\rightarrow 8)}$.
@@ IV-C Replicability and Verifiability
Our work will adhere to the Open Science principles[33]. We will prepare the replication package, containing raw data and scripts, and share it on Zenodo, according to FAIR principles [34], i.e., Findable, Accessible, Interoperable, and Reusable to foster applicability and verifiability and bolster the feasibility of future works.
@@ V Threats to Validity
In this section, we discuss the threats to the validity of our study. We categorized the threats into Construct, Internal, External, and Conclusion, following the guidelines defined by Wohlin et al. [70].
Construct Validity. Our specific design choices, including our measurement process and data filtering, may impact our results. To address this threat, we based our choice on past studies [19, 65, 21, 72] and used well-established guidelines in designing our methodology [57, 7].
Internal Validity. Our study relies on a large-scale analysis of 6,323 repositories, which can potentially be biased from the project selection. We address this threat by designing our inclusion criteria based on past studies [21, 72, 19].
@@ Acknowledgments
This work has been funded by the Research Council of Finland (grants n. 359861 and 349488 - MuFAno) and Business Finland (grant 6GSoft[4]).